Hi could you please show me how to do this? Thank you :)7.6 points Use the method of Lagrange multipliers to find the minimum value of f(x,y) = x2 + y2 on the line 2 + y = 12.
Answer #1

solution: given that flasy) = 21²+42 -o and subject to nery=120 let gensyn = x+y=12 Consider function! F(, 2) = f (27) 21, 2)